A publisher reports that 46% of their readers own a particular make of car. A marketing executive wants to test the claim that the percentage is actually different from the reported percentage. A random sample of 100 found that 38% of the readers owned a particular make of car. Determine the P-value of the test statistic. Round your answer to four decimal places.
Answer #1

Solution :

This is the two tailed test .

The null and alternative hypothesis is

H0 : p =0.46

Ha : p 0.46

= 0.38

P0 = 0.46

1 - P0 = 1 - 0.46=0.54

Test statistic = z

= - P0 / [P0 * (1 - P0 ) / n]

=0.38 -0.46 / [(0.46*0.54) / 100]

Z= −1.61

P(z < −1.61) = 0.0537 ( using z table )

P-value = 0.1074